# Chat2Workflow: A Benchmark for Generating Executable Visual Workflows with Natural Language

### 21 Apr 2026 v1

@@ Abstract
At present, executable visual workflows have emerged as a mainstream paradigm in real-world industrial deployments, offering strong reliability and controllability. However, in current practice, such workflows are almost entirely constructed through manual engineering: developers must carefully design workflows, write prompts for each step, and repeatedly revise the logic as requirements evolve—making development costly, time-consuming, and error-prone. To study whether large language models can automate this multi-round interaction process, we introduce Chat2Workflow, a benchmark for generating executable visual workflows directly from natural language, and propose a robust agentic framework to mitigate recurrent execution errors. Chat2Workflow is built from a large collection of real-world business workflows, with each instance designed so that the generated workflow can be transformed and directly deployed to practical workflow platforms such as Dify and Coze. Experimental results show that while state-of-the-art language models can often capture high-level intent, they struggle to generate correct, stable, and executable workflows, especially under complex or changing requirements. Although our agentic framework yields up to 5.34% resolve rate gains, the remaining real-world gap positions Chat2Workflow as a foundation for advancing industrial-grade automation.
@@
Yi Zhong♠♡, Buqiang Xu, Yijun Wang, Zifei Shan, Shuofei Qiao, Guozhou Zheng11footnotemark: 1, Ningyu Zhangthanks: Corresponding Author. Zhejiang University  Tencent {zhongyi0212, zhangningyu}@zju.edu.cn https://github.com/zjunlp/Chat2Workflow
@@ 1 Introduction
To study this problem systematically, we introduce Chat2Workflow, the first benchmark focused on generating executable visual workflows from natural language (Figure 1). Chat2Workflow is a high-quality dataset comprising 273 instances in six domains (Figure 2), to evaluate whether generated workflows meet the expected requirements for execution. Each round is equipped with realistic user requests and target workflow components across diverse domains and complexity levels.
The benchmark evaluates whether models can correctly infer workflow structure, select appropriate tools, and produce executable workflows that align with user intent. Experiments on Chat2Workflow show that current state-of-the-art language models can often capture high-level intent but struggle to generate correct and stable workflows, especially as tasks become more complex or requirements evolve. These results indicate that executable workflow generation remains a key bottleneck for practical system automation. Chat2Workflow provides a concrete benchmark for this challenge and points toward future research on structured reasoning and adaptive workflow synthesis for reliable systems.
@@ 2.1 Workflow Orchestration Platforms
Figure 3: Overview of Chat2Workflow benchmark construction and evaluation framework. Left: We collect workflows from six task domains (Research, Document, Enterprise, Developer, Education, AIGC) and reverse-engineer multi-turn instructions. Center: Users interact with LLMs through dialogue, and the model generates workflows in JSON format with CoT reasoning (node selection, design principles, and structured workflow). Right: The JSON workflow is converted to YAML format, uploaded to the platforms for execution, and evaluated against test cases to compute Pass Rate and Resolve Rate.
@@ 2.2 Task Formulation
Directly generating executable workflows is challenging for LLMs. To address this, we use a Chain-of-Thought (CoT) approach that outputs a simplified JSON. The model generates selected nodes $\hat{\mathcal{V}}$, planning rationale $p$, and the JSON representation $r$, wrapped in <node_selection>, <design_principle> and <workflow> tags. This JSON is then converted into an executable YAML file. The process is formalized as:
@@ 2.3 Dataset Construction
Figure 3 outlines our construction and evaluation framework. To address the absence of multi-turn workflow datasets, we collected production-level Dify and Coze configurations from official sources and GitHub. Based on these workflows derived from actual production, we reverse-engineered the workflow creation instruction set. To ensure that each round’s instruction is based on the previous ones and may add, modify, or refine workflow behavior, we selectively cluster the workflows in the same task context, then rewrite each of them as a round and the whole as a task. To facilitate subsequent evaluation, the written instructions clearly define the input and output variables and give explicit hints about the use of some key nodes. Ultimately, we obtained 27 tasks as the dataset, with each containing 2 to 4 rounds of instructions. 27 tasks can be further classified into the following six types according to the scenarios: Research, Document, Enterprise, Developer, Education and AIGC. More details are in Appendix A.
Although the same task objective can be achieved through different workflows, some specific nodes will inevitably be used, which will be recorded in advance for the corresponding rounds. After the collection and arrangement of the task instructions, we equipped each instruction with three test cases to actually carry out end-to-end testing. For task rounds with relatively clear outputs, the test cases will provide reference answers. For other general cases, only task input is provided. Test cases are primarily derived from existing public datasets, followed by scattered materials from the Internet, and finally synthesized data of language models. The design of test cases takes into account the task requirements of historical rounds, striving to comprehensively cover various situations of multi-branch workflows.
2.4 Data Selection and Processing
We selected the 20 most frequently used nodes to fully cover our dataset. To reduce generation complexity, we simplified the nodes’ I/O variable interfaces: the language agent only needs to provide key-value pairs for essential primary variables, while secondary variables default. In order to endow the language agent with the usage of each node and corresponding I/O interfaces, we craft node knowledge base for 20 node types, serving as a crucial part of the system prompt used for understanding the node ecosystem. On the basis of the above principles, we collect the official node documents, select the specified contents as needed for necessary rewriting, and then organize them into a predefined structured form.
@@ 2.5 Quality Control and Dataset Statistics
For the constructed data, we conduct a manual verification process for every individual input–output pair, ensuring the correctness, completeness, privacy and overall quality of the dataset. During this process, workflows that failed to meet our quality standards or violated predefined rules and compliance criteria are systematically identified and removed. This included cases with ambiguous instructions, invalid intermediate steps, time-varying sources, or incoherent workflow behaviors. After this rigorous filtering and validation process, we curate Chat2Workflow, a high-quality dataset consisting of 27 tasks with 79 multi-turn instructions in total, each accompanied by three test instances (totalling 237 instances) to assess whether the generated workflows adhere to the specified execution requirements.
@@ 2.6 Evaluation Protocol
Unlike general generation tasks such as text and images, the quality of a workflow cannot be judged entirely based on its representation form itself. It is essential to check whether it can be executed normally and what the execution outcome is. Therefore, we adopted a two-stage progressive evaluation method. The first stage aims at the workflow representation form itself, checking whether it is consistent with CoT sequence and whether the generated workflow JSON can be converted into a valid YAML file. The second stage involves verifying whether the workflow can be successfully imported into the platforms, correctly executed, and produce results that satisfy the specified requirements. We introduce two metrics, pass rate and resolve rate, corresponding to the first and second stages respectively.
@@
Pass rate measures the proportion of subtasks whose generated JSON representation is consistent with the CoT sequence as well as the predefined reference variables, and can be successfully converted into valid YAML files, following the sequential checks outlined in Algorithm 1.
If the above four steps are successfully completed, the workflow will be marked as passed and imported into the platforms. It can be calculated through the following formula:
@@
where %Pas. indicates Pass rate, $N_{\text{success\_subtask}}$ indicates the number of successful subtasks, $N_{\text{total\_subtask}}$ indicates the total number of subtasks.
Data: Workflow $\mathcal{G}$, instructions $\mathcal{I}$,Test case $\mathcal{T}=\{\text{input},\text{ref\_output}\}$
@@
Resolve rate measures the ratio of test cases where generated workflow can produce the results that meet all specified requirements of instructions. Algorithm 2 details the two-step verification process. Similarly, we will conduct the checks in sequence according to the process. It can be calculated through the following formula:
$\text{\%Res.}=\frac{N_{\text{success\_case}}}{N_{\text{total\_case}}},$ (4)
where %Res. refers to Resolve rate, $N_{\text{success\_case}}$ indicates the number of test cases whose execution results are in line with the requirement, $N_{\text{total\_case}}$ indicates the total number of test cases.
We employ DeepSeek-V3 for evaluation. The corresponding prompts related to the pass and resolve stages are detailed in Figures 6 and 7 of the Appendix, respectively. To assess performance, we conducted independent sampling for both stages, drawing 500 samples for the pass stage and 1,282 samples for the resolve stage. The LLM’s assessments demonstrated exceptional alignment with human evaluation, achieving 100% agreement for the pass stage and 98.83% agreement for the resolve stage.
@@ 3.1 Experimental Setup
Closed-Sourced
GPT-5.1 48.89 45.19 53.33 42.22 50.00 33.33 20.83 12.50 36.36 27.27 57.41 55.56 47.26 39.38
GPT-5.2 68.89 60.74 66.67 55.55 86.11 58.33 70.83 45.83 75.76 62.63 48.15 45.68 67.51 54.71
Claude-Sonnet-4.5 68.89 62.97 75.56 64.44 61.11 33.33 75.00 33.33 90.91 76.77 62.97 49.38 71.31 54.57
Gemini-3-Pro-Preview 77.78 77.78 82.22 77.04 86.11 75.93 70.83 52.78 90.91 72.73 74.08 66.67 80.17 71.59
Open-Sourced
Qwen-3-8B 40.00 21.48 8.89 4.45 8.33 0.00 12.5 2.78 3.03 0.00 7.41 7.41 13.92 6.89
Qwen-3-14B 26.67 21.48 33.33 17.78 13.89 0.00 20.83 12.50 27.27 5.05 35.18 25.31 27.43 15.19
Qwen-3-32B 31.11 26.67 28.89 14.07 30.56 8.33 50.00 27.78 33.33 28.28 48.15 33.34 36.71 23.35
Qwen-3-235B-A22B 35.55 32.59 51.11 38.52 19.45 2.78 62.50 29.17 39.39 32.32 42.59 27.78 40.93 27.71
Qwen-3-Coder-480B-A35B-Instruct 53.33 39.26 57.78 40.00 25.00 1.85 25.00 0.00 45.45 26.26 37.03 32.71 42.19 26.44
GLM-4.6 62.22 60.00 60.00 45.93 63.89 24.07 45.83 6.94 72.73 29.29 64.82 56.79 62.45 41.49
GLM-4.7 80.00 60.00 62.22 54.81 72.22 47.22 41.67 29.17 78.79 65.66 70.37 65.43 69.20 55.98
Deepseek-V3.1 51.11 51.11 51.11 36.29 61.11 26.85 20.83 4.17 45.45 39.39 62.96 55.56 51.48 39.24
Deepseek-V3.2 51.11 43.70 46.67 36.30 27.78 4.63 16.67 4.17 42.42 29.29 57.41 53.09 43.46 32.49
Kimi-K2-Instruct 51.11 48.89 37.78 33.33 11.11 8.33 29.17 9.72 21.21 15.15 51.85 48.76 36.29 31.08
Kimi-K2-Thinking 55.56 50.37 66.67 52.59 52.78 25.93 62.50 33.33 27.27 19.19 72.22 61.11 57.81 43.46
Table 1: Experimental results across six task domains for 4 closed-source and 11 open-source models. Pass Rate (%Pas.) measures format correctness, while Resolve Rate (%Res.) measures actual problem-solving capability.
Figure 5: Bad case analysis for the StudyPlanner task. We compare outputs from three representative models: Kimi-K2-Instruct produces an invalid edge connection in the iteration node; GPT-5.2 generates a correct and executable workflow; GLM-4.6 outputs logically inconsistent node declarations. This illustrates common failure modes in workflow generation.
@@ 3.2 Main Results
Q1: What is the gap between the workflow pass rate and resolve rate? For each kind of task, we show both the pass rate %Pas. and resolve rate %Res., representing the format legality and practical usability of workflows respectively. It can be observed that the resolve rate of all models is lower than the pass rate, with the largest disparity observed in GLM-4.6, reaching an average difference of 20.96%. Although Kimi-K2-Instruct shows the lowest gap of 5.21%, both of its absolute values are at a low level compared to the highest visible resolve rate of 71.59%. In addition, in Developer scenario, it has also shown a significant decline. In terms of specific task scenarios, the gap of GLM-4.6 in Education task scenarios has reached an astonishing 43.44%. We conclude that a legal workflow representation that meets the format requirements is still far from successful execution. The former merely serves as a fundamental and necessary condition, but not a sufficient one for achieving the latter, which often has the phenomenon of inflated scores.
Q2: To what extent do existing LLM agents approach the capabilities of real workflow design experts? Since the seed dataset is collected from typical workflows commmonly used by people, together with the mechanism for creating executable workflows through interactive dialogues and the requirements for using files as input and output, our benchmark can to a large extent represent the real-world scenarios. Considering whether the workflow can solve the ultimate problem only, Gemini-3-Pro-Preview performs the best on average, with its absolute performance of 71.59%, which is still far from the level of a practical and deployable workflow planner. Even the best open-source model GLM-4.7 can achieve 55.98% in resolve rate, it still struggles significantly in the Developer scenario. Given the limited number of node types considered, real-world workflow creation is inherently more complex; consequently, Gemini-3-Pro-Preview is supposed to show inadequate performance on complex workflows involving a larger set of nodes.
Q3: How can we improve the workflow generation capabilities of LLM agents? 1) First, by examining the performance of models in the Qwen-3 series, we can evaluate the impact of model size on workflow generation. From the perspective of average indicator, the performance improvement has been stable and significant as the number of model parameters has increased from 8B to 32B and then to 235B. We can draw a conclusion that a larger model size is conducive to workflow generation. 2) Second, post-training is a commonly used method to enhance models in a specific domain. In this domain, it can indeed make the generated workflow more in line with format specifications. However, a high pass rate does not necessarily lead to a high resolve rate. For example, in the Enterprise scenario GLM-4.6 has a higher pass rate than GLM-4.7, with 45.83% against 41.67%. But in terms of the resolve rate, GLM-4.7 is actually higher, with 29.17% against 6.94% Therefore, if only post-training is used to enhance compliance with the format, it may not be able to effectively improve the problem-solving ability of the workflow. 3) Last but not least, it is observed that the performance of instruct model is worse than think model. With larger model size, Qwen-3-Coder-480B-A35B-Instruct has a lower resolve rate than Qwen-3-235B-A22B, with 26.44% against 27.71%. Similarly, as both belong to the Kimi-K2 series, Kimi-K2-Thinking consistently outperforms the instruct model in all task domains. So the thinking mechanism can effectively promote the workflow generation capability.

### 26 May 2026 v2

@@ Abstract
At present, executable visual workflows have emerged as a mainstream paradigm in real-world industrial deployments, offering strong reliability and controllability. However, in current practice, such workflows are almost entirely constructed through manual engineering: developers must carefully design workflows, write prompts for each step, and repeatedly revise the logic as requirements evolve—making development costly, time-consuming, and error-prone. To study whether large language models can automate this multi-round interaction process, we introduce Chat2Workflow, a benchmark for generating executable visual workflows directly from natural language, and propose a robust agentic baseline to improve performance. The benchmark is built from a large collection of real-world business workflows, with each instance designed so that the generated workflow can be transformed and directly deployed to practical workflow platforms such as Dify and Coze. Experimental results show that while state-of-the-art language models can often capture high-level intent, they struggle to generate correct, stable, and executable workflows, especially given complex and evolving requirements. Although our agentic baseline yields up to 6.05% resolve rate gains, the remaining real-world gap positions Chat2Workflow as a foundation for advancing industrial-grade automation.
@@
Yi Zhong♠♡thanks: Equal contribution.     Corresponding Author., Yijun Wang11footnotemark: 1, Buqiang Xu11footnotemark: 1, Zifei Shan11footnotemark: 1, Shuofei Qiao, Guozhou Zheng, Ningyu Zhang22footnotemark: 2 Zhejiang University  Tencent {zhongyi0212, zhangningyu}@zju.edu.cn https://github.com/zjunlp/Chat2Workflow
@@ 1 Introduction
To systematically study this, we introduce Chat2Workflow (MIT license), the first benchmark focused on generating executable visual workflows from natural language (Figure 1). It comprises 237 human-annotated instances across six domains (Figure 2) to evaluate whether workflows meet the expected requirements for execution. Each round is equipped with realistic user requests and target workflow components across diverse domains and complexity levels.
The benchmark evaluates a model’s ability to infer structures, select tools, and produce executable, intent-aligned workflows. Experiments show that current state-of-the-art language models can often capture high-level intent but struggle to generate correct and stable workflows, especially as tasks become more complex or requirements evolve. To address this, we propose an error-driven agentic baseline with auto-repair mechanisms. Although it improves absolute resolve rates by 6.05%, the remaining performance gap highlights executable workflow generation as a key automation bottleneck, positioning Chat2Workflow as a crucial benchmark for future research in structured reasoning and adaptive workflow synthesis.
@@ 2.1 Workflow Orchestration Platforms
Figure 3: Overview of Chat2Workflow benchmark construction and evaluation framework. Left: The process of constructing multi-turn instructions involves six sequential stages: Prune, Filter, Cluster, Abstract, Chain, and Instantiate. Center: Users interact with LLMs through dialogue, and the model generates workflows in JSON format with CoT reasoning (node selection, design principles, and structured workflow). Right: The workflows span six distinct task scenarios. The JSON workflow is converted to YAML format, uploaded to the platforms for execution, and evaluated against test cases to compute Pass Rate and Resolve Rate.
@@ 2.2 Task Formulation
Directly generating executable workflows is challenging for LLMs. To address this, we use a Chain-of-Thought (CoT) approach that outputs a specially designed simplified JSON. The model generates selected nodes $\hat{\mathcal{V}}$, planning rationale $p$, and the JSON representation $r$, wrapped in <node_selection>, <design_principle> and <workflow> tags. This JSON is then converted into an executable YAML file. It can be formalized as:
@@
2.3 Node Selection and Processing
We selected the 20 most frequently used nodes to capture major real-world scenarios. To reduce generation complexity, we simplified the nodes’ input/output (I/O) variable interfaces: the language agent only needs to provide key-value pairs for essential primary variables, while secondary variables default. In order to endow the language agent with the usage of each node and corresponding I/O interfaces, we craft node knowledge base for 20 node types, serving as a crucial part of the system prompt used for understanding the node ecosystem. Based on these principles, we extract and rewrite essential information from official node documents into a predefined structure comprising <type>, <description>, <parameter>, <referable_variables>, and <supplementary_information>.
@@ 2.4 Dataset Construction
Figure 3 outlines our construction and evaluation framework. To build a multi-turn workflow dataset, three independent CS undergraduate volunteers processed production-level Dify and Coze configurations collected from official sources and GitHub through a systematic pipeline:
Pruning and Filtering. We first substituted unsupported workflow nodes with functional equivalents or pruned unviable functional components. Workflows with poor testability were discarded.
Clustering and Abstracting. The remaining workflows were clustered into six task scenarios: Research, Document, Enterprise, Developer, Education, and AIGC (details in Appendix A). Within each cluster, workflows were abstracted into natural language representations of their core functions, which are closely tied to essential key nodes.
Chaining. To construct multi-turn tasks (2–4 rounds), volunteers manually chained related workflows, treating each as a single turn. Each round’s instruction sequentially adds, modifies, or deletes functions from the previous round while maintaining logical consistency. We maximized variation across graph topologies, I/O types, and file formats, strictly distinguishing between function modifications and complete overwrites. I/O variables are explicitly formatted in parentheses, and unavoidable key nodes are recorded in advance as hints.
Instantiating. To enable end-to-end evaluation, we instantiated each instruction with three test cases that comprehensively cover historical requirements and multi-branch logic. Test cases were sourced hierarchically: (1) relevant Hugging Face datasets adapted to our I/O interfaces, (2) real-world internet materials, and (3) LLM-synthesized data. Reference answers are provided for deterministic outputs, whereas open-ended general cases consist solely of the provided inputs.
@@ 2.5 Quality Control and Dataset Statistics
We implement a strict cross-validation review process, with three volunteers auditing each other’s data to ensure correctness, completeness and privacy. To pass, the instructions must allow the target workflow’s key nodes and core functions to be manually recreated without ambiguity, exactly matching the creator’s intent. Cases with invalid intermediate steps, time-varying sources, or incoherent behaviors are systematically removed.
Following this rigorous validation, we curated Chat2Workflow, a high-quality dataset consisting of 27 tasks with 79 multi-turn instructions. Accompanied by 237 test instances (three per instruction), it robustly assesses whether generated workflows adhere to complex execution requirements.
@@ 2.6 Evaluation Protocol
Unlike standard text or image generation, workflows must be evaluated not only on their representational format but also on their executability and outcomes. Thus, we adopt a two-stage progressive evaluation: pass rate assesses structural and logical validity, while resolve rate evaluates functional execution correctness.
@@
Pass rate measures the proportion of subtasks that successfully pass the sequential checks in Algorithm 1. Specifically, a subtask passes if its JSON representation aligns with the CoT sequence and predefined variables, and converts into an importable YAML file. It is defined as:
@@
where $N_{\text{success\_subtask}}$ and $N_{\text{total\_subtask}}$ denote the number of passing subtasks and the total number of subtasks, respectively.
Resolve rate measures the proportion of test cases where the generated workflow correctly executes and produces outputs satisfying all instruction requirements in Algorithm 2. It is defined as:
$\text{\%Res.}=\frac{N_{\text{success\_case}}}{N_{\text{total\_case}}},$ (4)
where $N_{\text{success\_case}}$ and $N_{\text{total\_case}}$ represent the number of successful test cases and the total number of test cases, respectively.
Data: Workflow $\mathcal{G}$, instructions $\mathcal{I}$, Test case $\mathcal{T}=\{input,ref\_output\}$
@@
Closed-Sourced
GPT-5.1 48.89 41.48 53.33 40.74 50.00 33.33 20.83 8.33 36.36 19.19 57.41 45.06 47.26 34.46
GPT-5.2 68.89 53.33 66.67 50.37 86.11 58.33 70.83 38.89 75.76 54.55 48.15 32.10 67.51 47.40
Claude-Sonnet-4.5 68.89 56.30 75.56 62.22 61.11 27.78 75.00 31.94 90.91 60.61 62.97 41.97 71.31 47.96
Gemini-3-Pro-Preview 77.78 62.96 82.22 69.63 86.11 64.82 70.83 44.44 90.91 59.60 74.08 54.32 80.17 60.20
Open-Sourced
Qwen-3-8B 40.00 21.48 8.89 4.45 8.33 0.00 12.5 1.39 3.03 0.00 7.41 6.79 13.92 6.61
Qwen-3-14B 26.67 18.52 33.33 14.08 13.89 0.00 20.83 11.11 27.27 0.00 35.18 22.22 27.43 12.38
Qwen-3-32B 31.11 22.22 28.89 13.33 30.56 5.55 50.00 15.28 33.33 23.23 48.15 29.63 36.71 19.13
Qwen-3-235B-A22B 35.55 32.59 51.11 34.07 19.45 1.85 62.50 20.83 39.39 29.29 42.59 23.46 40.93 24.47
Qwen-3-Coder-480B-A35B-Instruct 53.33 29.63 57.78 28.89 25.00 1.85 25.00 0.00 45.45 17.17 37.03 23.46 42.19 19.13
GLM-4.6 62.22 57.04 60.00 39.26 63.89 24.07 45.83 5.55 72.73 25.25 64.82 39.51 62.45 35.02
GLM-4.7 80.00 54.82 62.22 45.18 72.22 44.45 41.67 19.45 78.79 52.53 70.37 50.61 69.20 46.55
Deepseek-V3.1 51.11 42.22 51.11 34.07 61.11 24.07 20.83 2.78 45.45 32.32 62.96 40.12 51.48 32.07
Deepseek-V3.2 51.11 40.74 46.67 36.29 27.78 1.85 16.67 4.17 42.42 22.22 57.41 39.51 43.46 27.43
Kimi-K2-Instruct 51.11 40.74 37.78 27.40 11.11 8.33 29.17 5.56 21.21 13.13 51.85 35.19 36.29 24.61
Kimi-K2-Thinking 55.56 44.44 66.67 46.67 52.78 25.93 62.50 26.39 27.27 15.15 72.22 43.83 57.81 36.00
Table 1: Experimental results across six task domains for 4 closed-source and 11 open-source models. Pass Rate (%Pas.) measures format correctness, while Resolve Rate (%Res.) measures actual problem-solving capability. The best results within each model series are marked in bold.
We employ GPT-5.1 for evaluation, sampling 355 instances for the pass stage and 463 for the resolve stage. To ensure reliability, three independent volunteers reviewed the LLM judge’s reasoning and verdicts. Discrepancies were handled by requiring strict unanimity: the agreement rate is defined as the percentage of cases where all three evaluators independently confirmed that the LLM strictly followed the evaluation prompt’s constraints (format, content, logic, and instruction following). Under this strict protocol, the LLM achieved 97.18% agreement for the pass stage and 93.09% for the resolve stage. Cross-model validation with DeepSeek-V3 demonstrates consistent stability (96.62% and 91.79%). However, it prioritizes formal compliance over deep semantic accuracy, unlike the stringent GPT-5.1. The exact prompts are detailed in Appendix (Figures 8 and 9).
@@ 3.2 Main Results
Q1: What is the gap between the workflow pass rate and resolve rate? For each task, we report both the pass rate and resolve rate to evaluate format compliance and practical usability, respectively. All models exhibit a lower resolve rate than pass rate. GLM-4.6 shows the largest disparity, averaging 27.43% and surging to 47.48% in the Education scenario. Although Kimi-K2-Instruct maintains a narrow gap of 11.68%, its absolute performance remains poor compared to the highest resolve rate (60.20%), particularly dropping in the Developer scenario. These results demonstrate that a syntactically legal workflow does not guarantee successful execution. Pass rates are necessary but insufficient, often inflating performance evaluations.
Q2: To what extent do existing LLM agents approach the capabilities of real workflow design experts? Our benchmark largely represents real-world scenarios through its use of typical workflows, interactive creation mechanisms, and file I/O requirements. Evaluated purely on end-to-end task resolution, Gemini-3-Pro-Preview performs best with an absolute resolve rate of 60.20%. However, It remains far below the threshold for a reliable, deployable planner. Even the best open-source model, GLM-4.7, achieves only a 46.55% resolve rate and shows significant weakness in Developer tasks. Given our limited set of node types, real-world workflow creation is inherently more complex. Consequently, we anticipate that even leading models will struggle further when confronted with larger node vocabularies.
Q3: How do scale, alignment, and reasoning impact workflow generation performance? 1) Scale. Increasing model parameters directly enhances performance. The Qwen-3 series demonstrates stable and significant improvements as its size scales from 8B to 32B and ultimately to 235B. 2) Alignment. While post-training improves format compliance, it does not guarantee problem-solving efficacy. For instance, in the Developer scenario, GLM-4.6 achieves a higher pass rate than GLM-4.7 (45.83% vs. 41.67%) but a drastically lower resolve rate (5.55% vs. 19.45%). Thus, optimizing solely for formatting may not effectively boost actual workflow usability. 3) Reasoning. Models equipped with reasoning capabilities consistently outperform standard instruction-tuned variants. Despite its massive scale, Qwen-3-Coder-480B-A35B-Instruct underperforms the smaller Qwen-3-235B-A22B in resolve rate (19.13% vs. 24.47%). Similarly, Kimi-K2-Thinking surpasses its instruct counterpart across all task domains, proving that explicit reasoning mechanisms are crucial for effective and robust workflow generation.
Q4: What shall we do to enhance the workflow generation capability of LLM agents? Analyzing failures from both the pass stage (see Figure 5 for the error distribution) and the resolve stage, we categorize the underlying errors into four primary dimensions. 1) Format. Syntactic and parsing rule violations, such as unclosed JSON brackets, redundant code fences, and improper character escaping. 2) Graph. Topological and structural flaws, including forward references violating topological sorting, hallucinated node connections, isolated nodes, and invalid cross-boundary edges connecting the inner and outer canvases of an iteration node. 3) Semantics. Cognitive misalignment regarding operational rules and node functionalities, manifested as failures to adhere to specific node configurations, violations of environment constraints, incorrect node types, or incompatible variable references. 4) Consistency. Functional or logical contradictions within the generation. This includes misalignments between the workflow’s actual function and the user’s core intent, mismatches with requested I/O variables, and logical discrepancies among different components of the model response or between consecutive turns. Therefore, addressing these critical bottlenecks is key to advancing LLM agents’ workflow generation capabilities.